OpenStudy (anonymous):

The system shown has _____ solution(s). y = x + 1 2y - x = 2 I'll give a medal

OpenStudy (anonymous):

\(\Large \begin{cases} y=x+1\quad (1) \\ 2y-x=2\quad (2) \end{cases}\) instead of \(\Large (1)\) into \(\Large (2)\) we get: \(\Large 2(x+1)-x=2\) \(\Large 2x+2-x=2\) \(\Large x=0\Rightarrow y=0+1=1\)